# SFMNet: Sparse Focal Modulation for 3D Object Detection

### v1

@@ Abstract
We propose SFMNet, a novel 3D sparse detector that combines the efficiency of sparse convolutions with the ability to model long-range dependencies. While traditional sparse convolution techniques efficiently capture local structures, they struggle with modeling long-range relationships. However, capturing long-range dependencies is fundamental for 3D object detection. In contrast, transformers are designed to capture these long-range dependencies through attention mechanisms. But, they come with high computational costs, due to their quadratic query-key-value interactions. Furthermore, directly applying attention to non-empty voxels is inefficient due to the sparse nature of 3D scenes. Our SFMNet is built on a novel Sparse Focal Modulation (SFM) module, which integrates short- and long-range contexts with linear complexity by leveraging a new hierarchical sparse convolution design. This approach enables SFMNet to achieve high detection performance with improved efficiency, making it well-suited for large-scale LiDAR scenes. We show that our detector achieves state-of-the-art performance on autonomous driving datasets.
@@ 1 Introduction
3D object detection has become fundamental due to its potential applications in autonomous driving [17, 2, 41, 29, 3, 50], and robotics [44, 66]. Extensive research has explored various methods, including point-based [32, 33, 34, 35, 36, 38, 39, 54], range-based [13, 42, 45], and voxel-based approaches [63, 5, 10, 23, 37, 52, 55, 56, 61, 62, 7, 47, 59, 58, 40]. Point-based methods process raw point clouds but often suffer from high computational costs or reduced accuracy in large-scale scenes due to point sampling. Range-based methods project 3D points to apply 2D convolutions, making them computationally efficient. However, they often struggle with occlusions, object localization, and size regression errors, as they lack the expressive 3D features necessary for accurate detection.
@@
Figure 1: SFMNet in comparison with other architectures. Similar to sparse convolution-based detectors, e.g.,[56] (a), our SFMNet detector (c) comprises a multi-stage 3D backbone with sparse convolution blocks. In contrast, SFMNet is fully sparse, including the 2D backbone, and integrates a novel SFM module (light orange) into both the 3D and 2D backbones. Furthermore, unlike transformers, SFMNet uses sparse convolutions as the token mixer, which makes query modulation efficient. Moreover, unlike the single-stride network design used in transformer-based detectors (e.g., [47]), our SFMNet detector employs an encoder-based design.
@@
Voxel-based methods, which are the focus of this paper, convert the irregular point cloud into regular voxels. This allows the use of convolutions, specifically sparse convolutions, which operate efficiently in the presence of empty voxels [52, 20, 9]. However, while efficient at capturing local structures, methods built on submanifold sparse convolution [19] significantly reduce the receptive field, limiting the representation capacity and long-range dependencies that are crucial for deep neural networks [49]. In 3D detection, long-range dependencies refer to connections between voxels within the same object and its immediate surroundings rather than between distant objects. Since LiDAR data is inherently sparse and occlusions are common, relying solely on local features makes it nearly impossible to detect objects like pedestrians with only a few sampled points. However, contextual cues, such as a crosswalk, can significantly aid detection, as even a few points above it may suggest the presence of a pedestrian. Therefore, long-range dependencies are essential for capturing contextual information.
@@
To address these limitations, the use of transformers has been proposed [46, 14, 47, 65, 43, 8, 48]. With self-attention (SA) as their core mechanism, transformers can capture global relationships, increasing the detector representation capacity. Despite its benefits, the quadratic complexity of SA is a significant bottleneck, particularly in autonomous vehicle scenarios where LiDAR scenes span hundreds of meters and contain hundreds of thousands of points.
@@
On the other hand, convolutional neural networks have recently made a comeback, primarily in 2D, through the use of large kernels [12, 11, 25]. Large-kernel convolution-based methods improve representation capacity by increasing the receptive field and offering greater robustness compared to transformers [4]. However, this approach is often impractical for 3D scenes, as the number of parameters grows cubically, necessitating techniques like sophisticated kernel weight partitioning with shared weights [6], data-driven kernel generation [27], or weight pruning [16]. Although these methods reduce parameter bottlenecks and enlarge the effective receptive field, they offer minimal accuracy improvements and are less efficient than traditional sparse convolutions.
@@
The Sparse Focal Modulation (SFM) module modulates short- and long-range contexts and integrates them with query tokens, similar to transformers. However, instead of the quadratic complexity of self-attention, SFM operates with linear complexity relative to the number of tokens (voxels). The core idea is to efficiently extract local contexts using sparse convolutions in a hierarchical structure, enabling the design to capture global relationships. Limiting attention to local context windows has been shown to reduce the complexity bottleneck [26]. However, due to the sparse nature of LiDAR points, the number of occupied voxels in each local window can vary significantly, making efficient local-based attention non-trivial [47]. To efficiently extract context windows, we design our SFM module with stacked submanifold sparse convolutions, using small kernels and varying dilation rates. This design enables hierarchical context aggregation, progressively expanding the receptive field at each level. Sparse convolutions capture local structures efficiently, while the hierarchical design and increased dilation rates allow for learning sparse long-range dependencies.
We validate the effectiveness of our detector on three well-known autonomous driving datasets. Our results demonstrate that the network advances the state of the art in long-range detection (Argoverse2 [50]), for which it is designed. On mid- to short-range datasets, namely Waymo Open [41] and nuScenes [2], it performs comparably or slightly better than previous state-of-the-art detectors.
@@
  • We propose a novel sparse focal modulation (SFM) module. It enables long-term dependencies, similar to transformers, while maintaining efficiency like sparse convolutions. This is achieved by hierarchically aggregating contextual information from sparse inputs. The SFM module outperforms traditional attention-based methods by reducing computational overhead and efficiently processing sparse voxels of varying densities within context windows.
@@ 2 Related work
3D detection on point clouds. 3D detection methods can be broadly categorized into point-based methods and voxel-based methods. Point-based methods use PointNet [32, 33]-like architectures to extract geometric features directly from raw point clouds [32, 33, 34, 35, 36, 38, 39, 54]. However, in large-scale point cloud scenes typical of outdoor LiDAR datasets in autonomous driving, these methods suffer from lower accuracy and high processing times due to point sampling and neighbor searching. In contrast, voxel-based methods convert irregular point clouds into regular voxel grids and apply sparse convolutions on non-empty voxels [63, 5, 10, 23, 37, 52, 55, 56, 61, 62, 7, 59, 58, 40]. While efficient for processing large-scale data, these methods rely on submanifold sparse convolution, which might limit network representation capacity by prioritizing efficiency over capturing long-range dependencies.
@@
Long-range dependencies for 3D detection. Recent 3D detection methods capture long-range dependencies using large-kernel sparse convolutions or transformers. Large-kernel convolution approaches expand the receptive field by increasing kernel sizes, a technique that has shown accuracy gains in 2D detection [12, 11, 25]. However, directly applying large kernels to 3D scenes incurs a cubically increasing computational overhead. To address this, LargeKernel3D [6] introduces a spatial partitioning convolution with shared weights among neighboring voxels. LinK [27] further optimizes efficiency by generating sparse kernels dynamically, assigning weights only to non-empty voxels rather than maintaining a dense kernel matrix. Similarly, LSK3DNet [16] reduces computation by pruning redundant weights across spatial and channel dimensions. While these methods alleviate some computational demands, they still face performance challenges in large-scale 3D data.
@@
Transformer-based detectors, by contrast, use self-attention to capture long-range dependencies, enabling global token interactions. However, applying attention in high-resolution 3D scenes with many tokens is computationally prohibitive [65, 8, 48]. To mitigate this, VoTr [30] implements local and dilated attention, focusing on sparse voxel interactions. SWFormer [43] utilizes a window-based transformer with a bucketing scheme to manage windows with varying numbers of non-empty voxels. SST [14] refines token processing by grouping tokens regionally, applying attention within each group, and shifting regions to capture structured interactions. DSVT [47] sequentially applies windowed self-attention along the X and Y axes to capture dependencies with reduced complexity. Although they improve long-range relationships, they still face quadratic scaling issues with context window tokens due to self-attention.
@@ 3 Method
Voxel-based detectors are widely used for 3D detection in outdoor scenes, such as those in autonomous driving scenarios. They commonly leverage sparse convolutions for efficient learning in sparse environments. There are two main types of sparse convolutions: regular sparse convolution [18] and submanifold sparse convolution [19]. Regular sparse convolutions expand features into (not necessarily occupied) neighboring voxels, enabling information exchange between spatially disconnected regions, but increasing feature map density. Due to their expansive nature, they are typically used for downsampling and upsampling. Submanifold sparse convolutions, on the other hand, operate strictly on occupied voxel features, ensuring that the output retains the same sparsity as the input, thereby facilitating efficient learning of local structures. However, due to their limited kernel size, they lack long-range dependencies, which are useful for object detection [49].
We aim to design an efficient detector built from sparse convolutions capable of capturing both short- and long-range dependencies. A straightforward approach to modeling global relationships would be to use transformers, which have shown great success in the 2D domain. However, applying standard transformers to sparse point clouds is challenging: (1) With numerous occupied voxels in each scene, standard self-attention suffers from massive query-key and query-value interactions, leading to a high computational overhead or suboptimal accuracy due to subsampling [31, 60]. (2) Due to the sparse nature of LiDAR point clouds, the number of occupied voxels in each local region can vary significantly, leading to inefficient batch processing [14, 47].
@@
The context extraction and interaction phases of our proposed SFM module are illustrated in Fig. 2, in comparison to naive local-based attention. Local-based attention involves a context extraction phase for each query token, followed by query-key and query-value interactions. The context extraction phase requires identifying all non-empty voxels within the context window, sampling or padding them to a fixed maximum number of neighbors, and concatenating their features for batch processing. In high-resolution scenes with hundreds of thousands of voxels, this process quickly becomes a bottleneck. Moreover, the query-key-value interactions result in quadratic complexity relative to the number of tokens within the context window. In contrast, SFM modulates a query’s interactions by extracting multiple levels of context using sparse convolutions, focusing on short-range structures at lower levels and long-range dependencies at higher levels. The sparse convolutions are designed to handle varying numbers of non-empty voxels efficiently within different context windows, enabling effective batch processing. By using (sparse) convolutions with different receptive fields for context extraction, the interaction phase is reduced to only the number of extracted context levels. Our SFM module is described in Section 3.1.
@@ 3.1 Sparse focal modulation
Focal Modulation (FM) is an alternative token mixer to self-attention, designed to limit token interactions. It has been applied in 2D tasks, such as classification, segmentation, and detection [53, 28]. Extending FM to 3D faces two primary challenges. First, 3D representations are inherently sparse. The voxel count, including both empty and non-empty voxels, scales cubically. For example, in Argoverse2 [50], a typical scene covers an area of $400\times 400\times 8$ meters, resulting in approximately $640$ million voxels at a resolution of $0.1\times 0.1\times 0.2$ meters. With only about 100,000 sampled points, this results in a high percentage of empty voxels. The second challenge is that 2D FM typically relies on large kernels to capture global context by using a stack of depth-wise convolutional layers with progressively larger kernels to capture visual contexts at various scales. However, in 3D object detection, sparse depth-wise convolutions have shown no advantage over spatial convolutions and are memory-intensive [6]. Furthermore, applying large kernels in 3D scenes leads to a cubic increase in parameter count, substantially raising computational costs.
We propose a new Sparse Focal Modulation (SFM) module that addresses these challenges. It captures both short- and long-range contexts while adaptively exchanging information with each token query. Unlike the conventional approach of using large kernels to expand the context window densely, we expand the context window sparsely with small kernels and increased dilation rates. We reason that queries related to small objects often rely on local structures and benefit from compact, focused context windows. In contrast, queries for large objects (e.g., partially occluded vehicles or roads) require broader context windows, but not necessarily dense windows, given the inherent sparsity of LiDAR point clouds. Hence, our context aggregation is sparse.
Fig. 3 illustrates our SFM module, which modulates each non-empty voxel feature ($x$) with surrounding information to produce a spatially aware feature ($z$). Its core component, the Context Aggregation (CA) block, captures relevant context for each query token by: 1) efficiently extracting contexts hierarchically, from local to global regions (e.g., red, green, and blue regions in Fig. 2); and 2) adaptively aggregating context features at each level using a gating mechanism to provide meaningful context for each query. The resulting aggregated context of the CA block, ctx, is modulated with the corresponding query $q$ to generate the feature $z$.
@@
Figure 3: Sparse Focal Modulation (SFM). Given non-empty voxel features $x$, the query $q$ and initial focal features $f^{0}$ are generated through linear layers. $f^{0}$ is fed into the Context Aggregation module, which aggregates multi-level contexts by stacking submanifold sparse convolutions (SubMConv) with small kernel sizes ($k$) and varying dilation rates ($d$), each followed by a GeLU activation. Context aggregation is achieved through a gated mechanism that adaptively combines features from each level ($f^{\ell}$) based on their relevance. The final context (ctx) is mapped to the query space using a linear layer $h(\cdot)$ and then modulated with the query $q$ to produce spatially-aware features $z$. This design captures both local and global dependencies while maintaining computational efficiency.
@@
where $\text{SubMConv}(\cdot;k^{\ell},d^{\ell})$ denotes a submanifold sparse convolution with kernel size $k^{\ell}$ and dilation rate $d^{\ell}$, and GeLU is an activation function [21]. The hierarchical convolutions effectively capture multiple levels of focal abstraction, with the effective receptive field at level $\ell$ given by:
@@
As for the second key principle, when aggregating context features, we note that not all queries benefit equally from short-range or long-range contexts. Specifically, queries for small objects are more likely to gain from the short-range context available at lower focal levels, while larger objects or background regions tend to benefit more from the long-range context captured at higher focal levels. To address this, a gating mechanism is employed to weigh the focal features from the different levels. This gate adaptively regulates the amount of context extracted from each level, enabling the CA module to learn suitable contexts for various queries. By combining the focal features with the gate weights, the CA produces a context feature vector for each query, defined as:
@@
Finally, we modulate the focal contexts to the queries through element-wise multiplication:
@@
This enables each query to learn its relevant context.
In summary, our module addresses the aforementioned challenges. By stacking sparse convolutions with small kernels and progressively increasing dilation, we effectively aggregate multi-level contexts. Efficiently handling voxel sparsity while modulating a query with its aggregated context maintains token interactions at linear complexity.
Method mAPH Vehicle Pedestrian Cyclist
Baseline 66.9 64.4 67.0 69.1
SFMNet 69.5 68.1 69.2 71.1
Table 1: The utility of our SFM module. Incorporating a single SFM module into each backbone (3D and 2D) within CenterPoint improves L2 mAPH by $2.6\%$ on a $20\%$ Waymo Open subset.
To demonstrate the utility of our proposed module, we incorporated a single SFM module into each backbone (3D and 2D) within the CenterPoint [56] architecture, without otherwise altering the structure. As shown in Table 1, this addition yielded a $2.6\%$ improvement in the Level-2 mAPH on a subset of the Waymo Open dataset [41]. Next, we will introduce our architecture that leverages our module.

### v2

@@ Abstract
We propose SFMNet, a novel 3D sparse detector that combines the efficiency of sparse convolutions with the ability to model long-range dependencies. While traditional sparse convolution techniques efficiently capture local structures, they struggle with modeling long-range relationships. However, these relationships are essential for 3D object detection. In contrast, transformers are designed to capture these long-range dependencies through attention mechanisms. But, they come with high computational costs, due to their quadratic query-key-value interactions. Furthermore, directly applying attention to non-empty voxels is inefficient due to the sparse nature of 3D scenes. Our SFMNet is built on a novel Sparse Focal Modulation (SFM) module, which integrates short- and long-range contexts with linear complexity by leveraging a new hierarchical sparse convolution design. This approach enables SFMNet to achieve high detection performance with improved efficiency, making it well-suited for large-scale LiDAR scenes. We show that our detector achieves state-of-the-art performance on autonomous driving datasets.
@@ 1 Introduction
3D object detection is fundamental for applications in autonomous driving [nuscenes, chang2019argoverse, kitti1, waymo, once, wilson2023argoverse] and robotics [thrun2006stanley, zhu2017target]. Extensive research has explored various methods, including point-based [qi2017pointnet, qi2017pointnet++, qi2018frustum, pointrcnn, pvrcnn, pvrcnn++, shi2020point, yang20203dssd], range-based [fan2021rangedet, sun2021rsn, tian2022fully], and voxel-based approaches [chen2019fast, chen2023voxelnext, voxelrcnn, pointpillars, parta2, shrout2023gravos, wang2023dsvt, second, ye2020hvnet, yin2021center, zhang2024safdnet, zhang2024hednet, zheng2021cia, zheng2021se, zhou2018voxelnet, song2023vp]. Point-based methods process raw point clouds but often suffer from high computational costs or reduced accuracy in large-scale scenes due to point sampling. Range-based methods project 3D points to apply 2D convolutions, making them computationally efficient. However, they often struggle with occlusions, object localization, and size regression errors, as they lack the expressive 3D features necessary for accurate detection.
@@
Figure 1: SFMNet in comparison with other architectures. Similar to sparse convolution-based detectors, e.g.,[yin2021center] (a), our SFMNet detector (c) comprises a multi-stage 3D backbone with sparse convolution blocks. In contrast, SFMNet is fully sparse, including the 2D backbone, and integrates a novel SFM module (light orange) into both the 3D and 2D backbones. Furthermore, unlike transformers, SFMNet uses sparse convolutions as the token mixer, which makes query modulation efficient. Moreover, unlike the single-stride network design used in transformer-based detectors (e.g., [wang2023dsvt]), our SFMNet detector employs an encoder-based design.
@@
Voxel-based methods, which are the focus of this paper, transform irregular point clouds into regular voxel grids. This enables the use of convolutions, specifically sparse convolutions, that operate efficiently despite the presence of many empty voxels [choy20194d, sparseconv, second]. While effective at capturing local structures, methods based on submanifold sparse convolution [graham2017submanifold] substantially limit the receptive field, thereby reducing the model’s ability to capture long-range dependencies, which are critical for detection [wang2018non]. In 3D detection, long-range dependencies encompass relationships among voxels within an object as well as between the object and its immediate surroundings. Given the inherent sparsity and frequent occlusions in LiDAR data, relying only on local features makes detecting small objects, such as pedestrians, extremely difficult when only a few points are available. Contextual information, such as the presence of a crosswalk, can greatly aid detection, as even a few points above it may indicate a pedestrian’s presence, highlighting the importance of long-range dependencies.
@@
To address these limitations, the use of transformers has been proposed [chen2023focalformer3d, fan2022embracing, sun2022swformer, vaswani2017attention, wang2023dsvt, wang2023unitr, zhou2022centerformer]. With self-attention (SA) as their core mechanism, transformers can capture global relationships, increasing the detector representation capacity. Despite its benefits, the quadratic complexity of SA is a significant bottleneck, particularly in autonomous vehicle scenarios where LiDAR scenes span hundreds of meters and contain hundreds of thousands of points.
@@
We propose a ’compromise’ between the two approaches by using large kernels in convolutional neural networks. This strategy has demonstrated efficiency, primarily in 2D [ding2022scaling, ding2024unireplknet, liu2022more]. Large-kernel methods improve representation capacity by increasing the receptive field and offering greater robustness compared to transformers [chen2024revealing]. However, this approach is often impractical for 3D scenes, as the number of parameters grows cubically, necessitating techniques like sophisticated kernel weight partitioning with shared weights [chen2023largekernel3d], data-driven kernel generation [lu2023link], or weight pruning [feng2024lsk3dnet]. Although these methods reduce parameter bottlenecks and enlarge the effective receptive field, they offer minimal accuracy improvements and are less efficient than traditional sparse convolutions.
@@
The Sparse Focal Modulation (SFM) module modulates short- and long-range contexts and integrates them with query tokens, similar to transformers. However, instead of the quadratic complexity of self-attention, SFM operates with linear complexity relative to the number of tokens (voxels). The core idea is to efficiently extract local contexts using sparse convolutions in a hierarchical structure, enabling the design to capture global relationships. Limiting attention to local context windows has been shown to reduce the complexity bottleneck [liu2021swin]. However, due to the sparse nature of LiDAR points, the number of occupied voxels in each local window can vary significantly, making efficient local-based attention non-trivial [wang2023dsvt]. To efficiently extract context windows, we design our SFM module with stacked submanifold sparse convolutions, using small kernels and varying dilation rates. This design enables hierarchical context aggregation, progressively expanding the receptive field at each level. Sparse convolutions capture local structures efficiently, while the hierarchical design and increased dilations allow for learning sparse long-range dependencies.
We validate the effectiveness of our detector on three well-known autonomous driving datasets. Our results demonstrate that the network advances the state-of-the-art in long-range detection (Argoverse2 [wilson2023argoverse]), for which it is designed. On mid- to short-range datasets, Waymo Open [waymo] and nuScenes [nuscenes], it performs comparably or slightly better than previous state-of-the-art detectors.
@@
  • We propose a novel sparse focal modulation (SFM) module that enables long-range dependencies like transformers while maintaining the efficiency of sparse convolutions. It hierarchically aggregates contextual information from sparse inputs, outperforming traditional attention methods by reducing computational overhead and efficiently handling sparse voxels of varying densities.
@@ 2 Related work
3D detection on point clouds. 3D detection methods are broadly categorized into point-based methods and voxel-based methods. Point-based methods use PointNet [qi2017pointnet, qi2017pointnet++]-like architectures to extract geometric features directly from raw point clouds [qi2017pointnet, qi2017pointnet++, qi2018frustum, pointrcnn, pvrcnn, pvrcnn++, shi2020point, yang20203dssd, wang2023sat]. However, in large-scale point cloud scenes typical of outdoor LiDAR datasets in autonomous driving, these methods suffer from lower accuracy and high processing times due to point sampling and neighbor searching. In contrast, voxel-based methods convert irregular point clouds into regular voxel grids and apply sparse convolutions on non-empty voxels [zhou2018voxelnet, chen2019fast, voxelrcnn, pointpillars, parta2, second, ye2020hvnet, yin2021center, zheng2021cia, zheng2021se, chen2023voxelnext, zhang2024hednet, zhang2024safdnet, shrout2023gravos, liu2024sparsedet]. While efficient for processing large-scale data, these methods rely on submanifold sparse convolution, which might limit network representation capacity by prioritizing efficiency over capturing long-range dependencies.
@@
Long-range dependencies for 3D detection. Recent 3D detection methods capture long-range dependencies using large-kernel sparse convolutions or transformers. Large-kernel convolution approaches expand the receptive field by increasing kernel sizes, a technique that has shown accuracy gains in 2D detection [ding2024unireplknet, ding2022scaling, liu2022more]. However, directly applying large kernels to 3D scenes incurs a cubically increasing computational overhead. To address this, LargeKernel3D [chen2023largekernel3d] introduces a spatial partitioning convolution with shared weights among neighboring voxels. LinK [lu2023link] further optimizes efficiency by generating sparse kernels dynamically, assigning weights only to non-empty voxels rather than maintaining a dense kernel matrix. Similarly, LSK3DNet [feng2024lsk3dnet] reduces computation by pruning redundant weights across spatial and channel dimensions. While these methods alleviate some computational demands, they still face performance challenges in large-scale 3D data.
@@
Transformer-based detectors, by contrast, use self-attention to capture long-range dependencies, enabling global token interactions. However, applying attention in high-resolution 3D scenes with many tokens is computationally prohibitive [zhou2022centerformer, chen2023focalformer3d, wang2023unitr]. To mitigate this, VoTr [mao2021voxel] implements local and dilated attention, focusing on sparse voxel interactions. SWFormer [sun2022swformer] utilizes a window-based transformer with a bucketing scheme to manage windows of varying non-empty voxels. SST [fan2022embracing] refines token processing by grouping tokens regionally, applying attention within each group, and shifting regions to capture structured interactions. DSVT [wang2023dsvt] sequentially applies windowed self-attention along the X and Y axes to capture dependencies with reduced complexity. Although they improve long-range relationships, they still face quadratic scaling issues with context-window tokens due to self-attention.
@@ 3 Method
Voxel-based detectors are widely used for 3D detection in outdoor scenes, such as those in autonomous driving scenarios. They commonly leverage sparse convolutions for efficient learning in sparse environments. There are two main types of sparse convolutions: regular [graham2015sparse] and submanifold [graham2017submanifold]. Regular sparse convolutions expand features into (not necessarily occupied) neighboring voxels, enabling information exchange between spatially disconnected regions, but increasing feature map density. Their expansive nature makes them suitable for downsampling and upsampling. Submanifold sparse convolutions operate only on occupied voxels, ensuring that the output retains the same sparsity as the input, thereby facilitating efficient learning of local structures. However, due to their limited kernel size, they lack long-range dependencies, which are useful for object detection [wang2018non].
We aim to design an efficient detector built from sparse convolutions capable of capturing both short- and long-range dependencies. While transformers are effective for modeling global relationships in 2D, applying them to sparse point clouds is challenging: (1) With numerous occupied voxels in each scene, standard self-attention suffers from massive query-key and query-value interactions, leading to a high computational overhead or suboptimal accuracy due to subsampling [pan20213d, zhao2021point]. (2) The varying voxel density across regions hinders efficient batching [fan2022embracing, wang2023dsvt].
@@
Our SFM module is described in Section 3.1. Fig. 2 illustrates its context extraction and interaction phases compared to naive local attention. Local attention involves extracting a context for each query token, then performing query-key and query-value interactions. This requires identifying all non-empty voxels in the context window, sampling or padding them to a fixed maximum, and concatenating features for batch processing. In high-resolution scenes with hundreds of thousands of voxels, this quickly becomes a bottleneck, and query-key-value interactions incur quadratic complexity with the number of tokens per window. In contrast, SFM modulates query interactions by extracting multiple context levels via sparse convolutions, capturing short-range structures at lower levels and long-range dependencies at higher levels. These sparse convolutions efficiently handle varying numbers of non-empty voxels within context windows, enabling effective batch processing. By using convolutions with different receptive fields for context extraction, the interaction phase is reduced to only the number of context levels.
@@ 3.1 Sparse focal modulation
Focal Modulation (FM) is an alternative token mixer to self-attention, designed to limit token interactions. It has been applied in 2D tasks like classification, segmentation, and detection [yang2022focal, ma2024efficient]. Extending FM to 3D faces two primary challenges. First, 3D representations are inherently sparse, with voxel count scaling cubically. For instance, a typical Argoverse2 [wilson2023argoverse] scene spans $400\times 400\times 8$ meters, yielding approximately $640$ million voxels at $0.1\times 0.1\times 0.2$ m resolution. With only about 100,000 sampled points, this results in a high percentage of empty voxels. Second, 2D FM typically captures global context using stacked depth-wise convolutions with progressively larger kernels to model multi-scale visual features. However, in 3D detection, sparse depth-wise convolutions offer no benefit over spatial ones and are memory-intensive [chen2023largekernel3d]. Furthermore, applying large kernels in 3D scenes leads to a cubic increase in parameter count, substantially raising computational costs.
We propose a new Sparse Focal Modulation (SFM) module that addresses these challenges and captures both short- and long-range contexts while adaptively exchanging information with each token query. Unlike the conventional approach of using large kernels to expand the context window densely, we expand the context window sparsely with small kernels and increased dilation rates. We reason that queries related to small objects often rely on local structures and benefit from compact, focused context windows. In contrast, queries for large objects (e.g., partially occluded vehicles or roads) require broader but not necessarily dense context windows, given the sparsity of LiDAR point clouds. Hence, our context aggregation is sparse.
@@
Figure 3: Sparse Focal Modulation (SFM). SFM leverages sparse convolutions to aggregate multi-level contexts and modulate them with query features, efficiently capturing both local and global dependencies.
Fig. 3 illustrates our SFM module, which modulates each non-empty voxel feature ($x$) with surrounding information to produce a spatially aware feature ($z$). Its core component, the Context Aggregation (CA) block, captures context for each query token by: 1) efficiently extracting contexts hierarchically, from local to global regions (e.g., red, green, and blue regions in Fig. 2); and 2) adaptively aggregating context features at each level using a gating mechanism to provide meaningful context for each query. The aggregated context of the CA block, ctx, is modulated with the corresponding query $q$ to generate the feature $z$.
@@
where $\text{SubMConv}(\cdot;k^{\ell},d^{\ell})$ denotes a submanifold sparse convolution with kernel size $k^{\ell}$ and dilation rate $d^{\ell}$, and GeLU is an activation function [hendrycks2016gaussian]. The hierarchical convolutions effectively capture multiple levels of focal abstraction, with the effective receptive field at level $\ell$ given by:
@@
When aggregating context features at each level, we note that not all queries benefit equally from short-range or long-range contexts. Specifically, queries for small objects are more likely to gain from the short-range context available at lower focal levels, while larger objects or background regions tend to benefit more from the long-range context captured at higher focal levels. To address this, a gating mechanism is employed to weigh the focal features from the different levels. This gate adaptively regulates the amount of context extracted from each level, enabling the CA module to learn suitable contexts for various queries. By combining the focal features with the gate weights, the CA produces a context feature vector for each query, defined as:
@@
Finally, we modulate the focal contexts (ctx) to the queries ($q$) through element-wise multiplication, which enables each query to learn its relevant context:
@@
In summary, to address voxel sparsity in 3D space, we propose Sparse Focal Modulation (SFM), which expands the context window efficiently using submanifold sparse convolutions with small kernels and increasing dilation. To handle varying contextual needs, the Context Aggregation (CA) block employs a gating mechanism to adaptively aggregate multi-scale features, allowing each query to focus on the most relevant context for accurate 3D detection.
To demonstrate its utility, we incorporated a single SFM module into each backbone (3D and 2D) within the CenterPoint [yin2021center] architecture, without otherwise altering the structure. This has already yielded a $2.6\%$ improvement in Level-2 mAPH on a $20\%$ subset of the Waymo Open Dataset [waymo].